Required information (The following information applies to the questions displayed below.] Santana Rey created Business Solut655 676 677 684 Advertising expense Mileage expense Miscellaneous expenses Repairs expense-Computer OOOO In response to requeHello if you could just make sure to do a-g id really appreciate it and I will thumbs up for your effort! thank you 9 The company received the balance due from Delta Co. for merchandise sold on February 23. 11 The company paid $900 cash for
Answer #1
Account Titles Debit Credit
a Computer Supplies Expense $              1,585 =730+2920-2065
       Computer Supplies $                1,585
(Computer supplies consumed)
b Insurance Expense $                  627
     Prepaid Insurance $                   627
(Prepaid insurance expired)
c Wages Expense $              1,225 =175*7
      Wages Payable     $                1,225
(Wages accrued but not paid)
d Rent Expense $              2,565
      Prepaid Rent $                2,565
(Prepaid rent expired)
e Depreciation Expense - Computer Equipment $              1,120
      Acc .Depreciation - Computer Equipment $                1,120
(Depreciation expense for the period)
f Depreciation Expense - Office Equipment $                  350
      Acc .Depreciation - Office Equipment $                   350
(Depreciation expense for the period)
g Cost of Goods Sold $              1,917 =7000-3480+620-7000*1%-476+9300-4610-9300*1%-2530-2092-1138-514
      Inventory $                1,917